A sample of Helium occupies 200 L at 30ºC under a pressure of 970 torr. What is the volume it occupies at standard temprature and pressure (STP)?

Grade:

1 Answers

Anurag Kishore
14 years ago

Hi, Cane

 

Use ideal gas eqn

P V/ T = constant

Here

P1  = 970 TORR     V1 = 200L      T1 = 30 + 273 = 303 K

P2 = 760 TORR ( AT STP)

V2 = x          T2 = 25 + 273 = 298 K

 

Now use

970 * 200/ 303 = 760 * x/ 298

=> x = 251.05 L

 

Thanks